Kevin started his science test at 2:34 p.m. and finished the test 75 minutes later. Kevin has football practice at 4:15 p.m. How much time did Kevin have between the end of the test and the beginning of football practice?
Answer by TimothyLamb(4379)

---
all times are after noon, so we can use noon as a reference:
---
start of test:
2:34 = 2*60 + 34 = 154 minutes (after noon)
---
end of test:
154 + 75 = 229 minutes (after noon)
---
start of football practice:
4:15 = 4*60 + 15 = 255 minutes (after noon)
---
time between end of test and start of football practice:
255 - 229 = 26 minutes
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---